The amount of loss (lower profit or higher cost) from not making the best decision for each state of nature is known as:

Choose one answer.

a. risk profile

b. opportunity loss

c. best payoff

d. utility

Question42

Marks: 1

A market research survey is available for $10,000. Using a decision tree analysis, it is found that the expected monetary value with no survey is $62,000. If the expected value of sample information is -$7,000, what is the expected monetary value with the survey?

Choose one answer.

a. $62,000

b. $45,000

c. $55,000

d. -$17,000

Question43

Marks: 1

Data shows that 75 out of 100 tourists who visit New York City visit the Empire State Building. It rains or snows in New York City one day in five. What is the joint probability that a randomly chosen tourist visits the Empire State Building on a day when it neither rains nor snows?

Choose one answer.

a. 0.95

b. 0.15

c. 0.60

d. 0.27

Question44

Marks: 1

The probability of a new Gaisano Mall being built in town is 64%. If Gaisano Mall comes to town, the probability of a new Davao Famous Restaurant being built is 90%. What is the probability of a new Gaisano Mall and a new Davao Famous Restaurant being built?

Choose one answer.

a. 0.675

b. 0.234

c. 0.576

d. 0.306

Question45

Marks: 1

A decision criterion which weights the payoff for each decision by its probability of occurrence is known as the:

Choose one answer.

a. probability

b. expected value criterion

c. expected value of perfect information

d. payoff criterion
Answer #1

41)

b. opportunity loss

42)

Expected monetory value with information = Expected value without information + Expected value of sample information

= 62,000 - 7000

c. $55,000

43)

Prob Empire = 0.75

Prob no rain = 4/5 = 0.8

Probability = 0.75 * 0.8 = 0.6

c. 0.60

44)

Prob = 0.64 * 0.9 = 0.576

c. 0.576

45)

b. expected value criterion

Please revert back in case of any doubt.

Please upvote. Thanks in advance.